Geometry. Find the perimeter of the triangle.
One side: 4
2nd side: √5 -√3
3rd side: √5 + √3

Answer by checkley75(3666)

4+SQRT5-SQRT3+SQRT5+SQRT3=4+2SQRT5=4+2*2.236=4+4.472=8.472